# Towards Personalized Deep Research: Benchmarks and Evaluations

### v1

@@ Abstract
Deep Research Agents (DRAs) can autonomously conduct complex investigations and generate comprehensive reports, demonstrating strong real-world potential. However, existing evaluations mostly rely on close-ended benchmarks, while open-ended deep research benchmarks remain scarce and typically neglect personalized scenarios. To bridge this gap, we introduce Personalized Deep Research Bench, the first benchmark for evaluating personalization in DRAs. It pairs 50 diverse research tasks across 10 domains with 25 authentic user profiles that combine structured persona attributes with dynamic real-world contexts, yielding 250 realistic user-task queries. To assess system performance, we propose the PQR Evaluation Framework, which jointly measures (P) Personalization Alignment, (Q) Content Quality, and (R) Factual Reliability. Our experiments on a range of systems highlight current capabilities and limitations in handling personalized deep research. This work establishes a rigorous foundation for developing and evaluating the next generation of truly personalized AI research assistants.
@@ 2.2 Benchmarking Personalization Performance
Meanwhile, most personalization benchmarks focus on general tasks and remain insufficient for complex deep research scenarios. LaMP (Salemi et al., 2024) introduces seven classification and generation tasks to evaluate the personalized output capacity of LLMs. PersonaGym (Samuel et al., 2025) introduces PersonaScore to evaluate the adherence of LLM agents to assigned personas at scale. PersonalLLM (Zollo et al., 2025) uses reward models to act as different user personas to evaluate response preference. AI Persona (Wang et al., 2024) concentrates on the lifelong learning of user profiles with LLM-as-a-judge evaluation. Additionally, PersonaMem (Jiang et al., 2025) benchmarks the adaptability of LLMs to evolving user personas. PersonaFeedback (Tao et al., 2025) provides a large human-annotated benchmark for response tailoring to explicit personas. PersonaLens (Zhao et al., 2025) introduces LLM-based user and judge agents to assess personalization and task success in realistic dialogues.
@@ 3.1 Deep Research Task Formulation
To begin, we defined a set of 10 distinct domains, $\mathcal{D}=\{d_{1},d_{2},...,d_{10}\}$ covering major and impactful aspects of daily life (e.g., Education, Career Development, Healthcare, Financial Planning). To ensure that the tasks within each domain are both realistic and practically relevant, we collaborated with a diverse group of domain experts, such as travel bloggers, financial advisors and educational consultants, to design the initial set of tasks.
@@
To complement these explicit personas with dynamic context, we employed professional annotators simulate the daily interactions of these collected personas through a phone APP. Over a period, they were instructed to: 1) Record naturalistic memory snippets ($m_{j}$), such as travel aspirations, health goals, and family plans; and 2) Conduct conversational interactions ($c_{j}$) with the basic assistant integrated in the app. This longitudinal data captures each user’s evolving interests, habits, and implicit preferences. These multi-modal data streams were then processed by the built-in management system of the APP, $f_{\theta}$, to generate dynamic personalized contexts: $\mathcal{P}_{c}=\{Pc_{j}\mid Pc_{j}=f_{\theta}(m_{j},c_{j}),\ j=1,\dots,25\}$. Annotation details are in Appendix H.
@@ 4.1 P - Personalization Alignment
Evaluating the personalization of generated report is a significant challenge due to its subjective, multi-dimensional nature. Traditional static metrics with uniform standard overlook the nuanced alignment between reports and user profile. To address this, we introduce Personalization Alignment (P-Score), a dynamic evaluation framework that generates customized criteria and scores for each user–task pair.
@@ 4.2 Q – Quality of Content
We assess report reliability via an automated factual grounding framework, inspired by ResearcherBench (Xu et al., 2025) and DeepResearch Bench (Du et al., 2025). The process has three stages:
@@
A Judge LLM is employed to extract all verifiable factual claims with their sources, forming a set of triplets $\mathcal{TRI}=\{(c_{i},\text{idx}_{i},\text{source}_{i})\}_{i=1}^{N}$, where uncited claims have empty sources. A second pass deduplicates claims:
@@ 4.4 Final Score Aggregation
To obtain a holistic measure of the report, we define the overall score as a generalized mean over the three scores:
$S_{\mathrm{overall}}=\frac{\sum_{a\in\{P,Q,R\}}\lambda_{a}S_{a}}{\sum_{a\in\{P,Q,R\}}\lambda_{a}}$ (6)
where $S_{a}\in\{S_{P},S_{Q},S_{R}\}$ and $\lambda_{a}$ are their weights (default $\lambda_{a}=1$). This aggregation provides a flexible and comprehensive measure of the personalized deep research report.
@@ 5.1 Experimental Settings
We benchmarked a diverse set of systems, including commercial deep research systems: Gemini-2.5-Pro Deep Research, O3 Deep Research, Perplexity Deep Research (Google DeepMind, 2025; OpenAI, 2025b; Perplexity Team, 2025), open-source deep research agents: Deerflow, Oagents, Miroflow (ByteDance, 2025a; Zhu et al., 2025a; MiroMind AI Team, 2025), and leading LLMs with search tools: Gemini-2.5-Pro-Search, Claude-3.7-Sonnet-Search, Perplexity-Sonar-Reasoning-Pro, GPT-4.1-Search-Preview (DeepMind, ; Anthropic, ; AI, 2025; OpenAI, a). Due to computational constraints, the evaluation was performed on a subset of 150 representative queries. GPT-5 (OpenAI, 2025a) was utilized as the judge model for Personalization (P) and Quality (Q) metrics, while the more efficient GPT-5-Mini (OpenAI, b) served as the judge for the Reliability (R) metric, ensuring a balance of advanced reasoning and efficiency (more details in Appendix B).
@@ 5.3 Impact of Information Availability on Personalization
While the previous section evaluated agents with explicit personas, a more realistic scenario involves inferring user needs from conversational or interaction context, as explicit personas are rarely available. To examine this, we conducted a comparative experiment under three conditions: Task Only (the agent receives only the task), Task w/Context (the task plus user’s conversational or interaction context), and Task w/Persona (the task plus an explicit user persona, consistent with our main experiment). We present the results in Table 2 and Figure 3. A more detailed results are in Appendix F.
@@ 5.5 Alignment with Human Consistency
Based on the results shown in Table 4, GPT-5 achieved the highest PCA and lowest MARD, indicating the strongest agreement with human judgments, while maintaining a reasonable cost per evaluation. We finally select GPT-5 as our primary judge model.
@@
GPT-5 0.43 1.40 0.32
Claude-3.7-Sonnet 0.39 1.44 0.60
Gemini-2.5-Pro 0.40 2.33 0.21
@@ Appendix D Persona Schema
Table 5: Predefined Schema For Persona Collection
@@ Appendix E Definitions of Evaluation Dimensions
Appendix F A More Detailed Experiments Results
@@
Appendix G Case Study
@@
Appendix H Data Annotation
@@
Appendix I Prompt Templates
I.1 Prompts in Personalization Evaluation
@@
I.2 Prompts in Quality Evaluation
@@
I.3 Prompts in Reliability Evaluation

### v2

@@ Abstract
Deep Research Agents (DRAs) can autonomously conduct complex investigations and generate comprehensive reports, demonstrating strong real-world potential. However, existing evaluations mostly rely on close-ended benchmarks, while open-ended deep research benchmarks remain scarce and typically neglect personalized scenarios. To bridge this gap, we introduce Personalized Deep Research Bench (PDR-Bench), the first benchmark for evaluating personalization in DRAs. It pairs 50 diverse research tasks across 10 domains with 25 authentic user profiles that combine structured persona attributes with dynamic real-world contexts, yielding 250 realistic user-task queries. To assess system performance, we propose the PQR Evaluation Framework, which jointly measures Personalization Alignment, Content Quality, and Factual Reliability. Our experiments on a range of systems highlight current capabilities and limitations in handling personalized deep research. This work establishes a rigorous foundation for developing and evaluating the next generation of truly personalized AI research assistants111  Code: https://github.com/OPPO-PersonalAI/PersonalizedDeepResearchBench..
@@ 2.2 Benchmarking Personalization Performance
Meanwhile, most personalization benchmarks focus on general tasks and remain insufficient for complex deep research scenarios. LaMP (Salemi et al., 2024) introduces seven classification and generation tasks to evaluate the personalized output capacity of LLMs. PersonaGym (Samuel et al., 2025) introduces PersonaScore to evaluate the adherence of LLM agents to assigned personas at scale. PersonalLLM (Zollo et al., 2025) uses reward models to act as different user personas to evaluate response preference. AI Persona (Wang et al., 2024b) concentrates on the lifelong learning of user profiles with LLM-as-a-judge evaluation. Additionally, PersonaMem (Jiang et al., 2025) benchmarks the adaptability of LLMs to evolving user personas. PersonaFeedback (Tao et al., 2025) provides a large human-annotated benchmark for response tailoring to explicit personas. PersonaLens (Zhao et al., 2025) introduces LLM-based user and judge agents to assess personalization and task success in realistic dialogues.
@@ 3.1 Deep Research Task Formulation
To begin, we defined a set of 10 distinct domains, $\mathcal{D}=\{d_{1},d_{2},...,d_{10}\}$ covering major and impactful aspects of daily life (e.g., Career Development, Education, Healthcare, Financial Planning). To ensure that the tasks within each domain are both realistic and practically relevant, we collaborated with a diverse group of domain experts, such as travel bloggers, financial advisors and educational consultants, to design the initial set of tasks.
@@
To complement these explicit personas with dynamic context, we employed professional annotators simulate the daily interactions of these collected personas through a phone APP. Over a period, they were instructed to: 1) Record naturalistic memory snippets ($m_{j}$), such as travel aspirations, health goals, and family plans; and 2) Conduct conversational interactions ($c_{j}$) with the intelligent assistant integrated in the app. This longitudinal data captures each user’s evolving interests, habits, and implicit preferences. These multi-modal data streams were then processed by the built-in management system of the APP, $f_{\theta}$, to generate dynamic personalized contexts: $\mathcal{P}_{c}=\{Pc_{j}\mid Pc_{j}=f_{\theta}(m_{j},c_{j}),\ j=1,\dots,25\}$. Annotation details are in Appendix I.
@@ 4.1 P - Personalization Alignment
Evaluating the personalization of generated report is a significant challenge due to its subjective, multi-dimensional nature. Recent studies (Wang et al., 2024a; Guan et al., 2025; Zhu et al., 2025c) consistently emphasize that personalization evaluation should move beyond global correctness toward preference-aware and user-centered assessment paradigms, highlighting the need for individualized criteria rather than generic evaluation frameworks. Motivated by these insights and our requirement analysis, we introduce Personalization Alignment (P-Score), a dynamic evaluation framework that generates customized criteria and scores for each user–task pair.
@@ 4.2 Q – Quality of Content
Although factuality metrics (Min et al., 2023; Wei et al., 2024; Chern et al., 2023) exist, they are primarily designed to verify atomic facts against knowledge bases or external search results. They remain unsuitable for deep research setting, where factuality must be evaluated through retrieved citations to assess both the factual reliability of the report and the agent’s capacity in utilizing web information. We therefore assess report reliability via an automated factual grounding framework, inspired by ResearcherBench (Xu et al., 2025) and DeepResearch Bench (Du et al., 2025). The process has three stages:
@@
A Judge LLM is employed to extract all verifiable factual claims with their sources, forming a set of triplets $\mathcal{TRI}=\{(c_{i},\text{idx}_{i},\text{source}_{i})\}_{i=1}^{N}$, where uncited claims have empty sources (see Appendix J.3). A second pass deduplicates claims:
@@ 4.4 Final Score Aggregation
To obtain a holistic measure of the report, we define the final overall score as a arithmetic mean over the three dimension scores:
$S_{\mathrm{overall}}=\frac{S_{P}+S_{Q}+S_{R}}{3}$ (6)
where $S_{P}$, $S_{Q}$, and $S_{R}$ denote the scores for personalization, quality and factual reliability respectively. This aggregation provides a straightforward and comprehensive measure of the personalized deep research report.
@@ 5.1 Experimental Settings
We benchmarked a diverse set of systems, including commercial deep research systems: Gemini-2.5-Pro Deep Research, O3 Deep Research, Perplexity Deep Research (Google DeepMind, 2025; OpenAI, 2025b; Perplexity Team, 2025), open-source deep research agents: Deerflow, Oagents, Miroflow (ByteDance, 2025a; Zhu et al., 2025a; MiroMind AI Team, 2025), and leading LLMs with search tools: Gemini-2.5-Pro-Search, Claude-3.7-Sonnet-Search, Perplexity-Sonar-Reasoning-Pro, GPT-4.1-Search-Preview (DeepMind, ; Anthropic, ; AI, 2025; OpenAI, a). Due to computational constraints, the evaluation was performed on a subset of 150 representative queries. GPT-5 (OpenAI, 2025a) was utilized as the judge model for Personalization (P) and Quality (Q) metrics, while the more efficient GPT-5-Mini (OpenAI, b) served as the judge for the Reliability (R) metric, ensuring a balance of advanced reasoning and efficiency (more details in Appendix B and  F).
@@ 5.3 Impact of Information Availability on Personalization
While the previous section evaluated agents with explicit personas, a more realistic scenario involves inferring user needs from conversational or interaction context, as explicit personas are rarely available. To examine this, we conducted a comparative experiment under three conditions: Task Only (the agent receives only the task), Task w/Context (the task plus user’s conversational or interaction context), and Task w/Persona (the task plus an explicit user persona, consistent with our main experiment). We present the results in Table 2 and Figure 3. A more detailed results are in Appendix G.
@@ 5.5 Alignment with Human Consistency
Based on the results shown in Table 4, GPT-5 achieved the highest PCA and lowest MARD, indicating the strongest agreement with human judgments, while maintaining a reasonable cost (\$0.68 per query). We finally select GPT-5 as our primary judge model.
@@
GPT-5 0.43 1.40 0.68
Claude-3.7-Sonnet 0.39 1.44 0.97
Gemini-2.5-Pro 0.40 2.33 0.61
@@ 7 Ethics Statement
8 Reproducibility statement
To ensure reproducibility, we provide comprehensive details across the paper and appendix. Section 3 and Appendix I describes benchmark construction, including task design, user profiles collection, and query pairing. The evaluation framework is detailed in Section 4, including the dynamic weight allocation, granular criterion generation, and scoring methodology for each dimension. Section 5 and Appendix B outline experimental setups, systems, and configurations. Prompt templates (Appendix J), persona schema (Appendix D), and evaluation dimensions (Appendix E) are fully documented. We have submitted our evaluation data in the Supplementary Material. Due to OpenReview’s file size limit, we only upload a subset of them. We will fully release the benchmark immediately after the double blind review process. We invite the community to build upon this work in advancing personalized deep research agents.
@@ Appendix D Persona Schema
Table 5: Predefined Schema For Persona Collection (All Private Data Anonymized for Release)
@@ Appendix E Definitions of Evaluation Dimensions
Appendix F Running Cost
Table 8: Running Cost Comparison Across Models and Personalization Settings (per query)
Model / Setting Task Only Task w/ Persona Task w/ Context
Open-Source DRAs (GPT-5-Mini based)
OAgents \$1.60 \$1.70 \$3.00
DeerFlow \$0.50 \$0.57 \$1.20
MiroFlow \$1.00 \$1.11 \$2.10
LLM with Search Tools
Gemini-2.5-Pro w/ Search \$0.05 \$0.06 \$0.24
Claude-3.7-Sonnet w/ Search \$0.03 \$0.04 \$0.31
Perplexity-Sonar-Reasoning-Pro \$0.02 \$0.03 \$0.78
GPT-4.1 w/ Search \$0.02 \$0.02 \$0.31
Appendix G A More Detailed Experiments Results
@@
Appendix H Case Study
@@
Appendix I Data Annotation
@@
Appendix J Prompt Templates
J.1 Prompts in Personalization Evaluation
@@
J.2 Prompts in Quality Evaluation
@@
J.3 Prompts in Reliability Evaluation